OpenStudy (anonymous):

A flashlight battery has a lifetime of 5 hours and it is rated for 0.6 ampere-hours. (a) What is the current that can be delivered by the battery? (b) How much power can be delivered if the terminal voltage is 8 V? (c) How much energy is stored in the battery?

OpenStudy (anonymous):

a. If it has .6 ampere-hours means the battery can run a current of x amps for .6/x hours From this we see that the current is .6/5hours = 0.12 amperes b. Power= I*V= .12A*8V= 0.96 Watts c. Power= energy/second Energy= Power*#seconds= 0.96 Watts*5hours*(3600seconds/hour)=17280 joules or E=17.28 kilojoules